### v1

# Prompt-Guided Latent Diffusion with Predictive Class Conditioning for 3D Prostate MRI Generation

Abstract
Latent diffusion models (LDM) could alleviate data scarcity challenges affecting machine learning development for medical imaging. However, medical LDM training typically relies on performance- or scientific accessibility-limiting strategies including a reliance on short-prompt text encoders, the reuse of non-medical LDMs, or a requirement for fine-tuning with large data volumes. We propose a Class-Conditioned Efficient Large Language model Adapter (CCELLA) to address these limitations. CCELLA is a novel dual-head conditioning approach that simultaneously conditions the LDM U-Net with non-medical large language model-encoded text features through cross-attention and with pathology classification through the timestep embedding. We also propose a joint loss function and a data-efficient LDM training framework. In combination, these strategies enable pathology-conditioned LDM training for high-quality medical image synthesis given limited data volume and human data annotation, improving LDM performance and scientific accessibility. Our method achieves a 3D FID score of 0.025 on a size-limited prostate MRI dataset, significantly outperforming a recent foundation model with FID 0.071. When training a classifier for prostate cancer prediction, adding synthetic images generated by our method to the training dataset improves classifier accuracy from 69% to 74%. Training a classifier solely on our method’s synthetic images achieved comparable performance to training on real images alone.
Keywords Generative AI  $\cdot$ Large Language Model Adapters  $\cdot$ Medical Image Generation  $\cdot$ Magnetic Resonance Imaging  $\cdot$ Prostate Cancer
1 Introduction
Machine learning (ML) for medical image interpretation is a rapidly expanding field intent on improving imaging-based healthcare. While ML continues to improve outside of medical imaging, recent ML advancements require training data in higher orders of magnitude than their predecessors. These advances cannot translate easily to the medical imaging space due to labeled data scarcity challenges including patient privacy and the cost of expert annotation. Generative approaches, including latent diffusion models (LDM), provide a possible solution to labeled training data scarcity, whereby synthetic data could be used to artificially increase ML training data volume [1].
Developing LDMs for medical imaging poses a few challenges. Training LDMs from scratch require large volumes of data. The Stable Diffusion LDM [2] for generating nonmedical, 2D color images was trained initially on 400 million image-text pairs [3], with subsequent versions training on 5.85 billion image-text pairs [4]. These data volumes are inaccessible to most healthcare institutions due to aforementioned data scarcity challenges. One approach is to fine-tune non-medical (2D) LDMs on 2D medical imaging data [5], however this strategy cannot easily extend to 3D. As an added challenge, because 3D medical images typically have more voxels than 2D images have pixels, 3D medical image generation should theoretically require more training data. Another approach is to add labeled data conditioning to increase data efficiency during training [6]. While some medical LDMs have attempted to use the pre-trained CLIP text encoder to condition the LDM, similar to Stable Diffusion [2], research has shown that the CLIP text encoder mostly emphasizes the first 20 tokens [7]. Thus, rather than using detail-rich clinical reports, medical imaging LDM studies typically use only short, curated medical phrases or keywords with either the existing [8] or fine-tuned [5] CLIP text encoder. These strategies require extra annotation and preprocessing or volumes of data that are challenging for a single institution to gather, limiting the scientific accessibility of medical LDM development.
In this study we propose Class-Conditioned Efficient Large Language Model Adapter (CCELLA), a text adapter extending ELLA [9], to improve LDM performance for synthetic medical image generation in a challenging medical imaging domain. CCELLA processes the text embedding generated by a large language model (LLM) from clinical reports prior to dual-head conditioning of the LDM U-Net. CCELLA performs two functions simultaneously: the first head aligns the text embedding with the image latent space prior to conditioning; the second head provides class conditioning for the LDM during training. Each of these functions have independently demonstrated improvements in LDM performance outside of medical imaging [9, 10]. This joint function is hypothesized to provide two advantages: 1) to assert that the LDM receives conditioning about the presence or absence of pathology within the image, 2) more importantly, to inform CCELLA which text features should be emphasized when aligning the embedding to the image latent space for LDM conditioning. To achieve these functions, we also propose a modified loss function that simultaneously trains the LDM to produce high quality medical images and to appropriately extract pathology information from the clinical report text.
Finally, we propose an LDM training framework around CCELLA that uses limited human data annotation and a training data volume reasonably accessible to a single healthcare institution, with the intent of improving the scientific accessibility of LDMs for medical imaging despite limited available data volumes. The framework functions by exploiting the clinical report and pathology labeling already performed by radiologists during clinical routine and by reducing training requirements with publicly available, pre-trained components [1, 11].
This study focused specifically on MRI for prostate cancer as the target domain to evaluate this framework. This pathology and imaging method was selected due to its difficulty in clinical interpretation, its moderate inter-reader variability between clinical experts, and the relatively low-availability of publicly available domain-specific data with the largest dataset consisting of 1500 samples [12, 13]. This study then performed a series of ablations to assess the mechanism by which CCELLA improves LDM performance and compared performance differences among variants of the proposed adapter. Finally, to validate utility of the proposed LDM training framework for dataset augmentation, this study demonstrated that the synthetic data generated with our proposed LDM framework can improve a classifier model at interpreting prostate MRI for cancer presence.
In summary, this paper proposes the following:
  • A novel dual-head adapter for simultaneous alignment of medical text features extracted by a non-medical LLM to the medical imaging space and pathology class extraction from medical text
  • A LDM training framework for improved pathology-conditioned synthetic medical image generation that is trainable with a reasonable data volume.
  • A method for augmenting datasets to improve downstream ML tasks using LDM-generated synthetic images with limited data annotation beyond clinical routine.
2 Related Work
2.1 Latent Diffusion Models
Latent diffusion models (LDM) are a type of generative model consisting of two stages: a vector-quantized variational autoencoder (VQ-VAE) and a denoising U-Net [2]. The VQ-VAE compresses the image to a lower-dimensional representation in a discretized latent space [14]. Functionally this reduces computational complexity increasing both training and inference speed, and coerces the image representation to fit a normal distribution. The denoising U-Net learns the reverse process to a Markov Chain [15]. During LDM training, the Markov Chain forward process gradually adds Gaussian noise to the latent image for each of many, e.g. 1000, timesteps. The denoising U-Net is then trained to reverse the noise addition process by predicting the noise added between the true latent image and a randomly selected timestep up to 1000.
At inference time, the synthetic latent image begins as randomly-generated Gaussian noise. The U-Net then proceeds to denoise the image from timestep 1000 to 0, at which point the resultant latent image is decoded by the VQ-VAE to yield a synthetic image.
2.2 Latent Diffusion Models within Medical Imaging
LDMs have shown promising results in generating synthetic images both outside and within medical imaging [16]. Because LDM research focuses primarily on 2D, non-medical color image generation on large datasets, medical imaging LDM studies typically also require large data volumes or strategic data-efficient methods. For example, Roentgen [5] adapted non-medical LDMs for chest x-ray generation by fine-tuning the Stable Diffusion LDM [2] on a public dataset with over 350k images. Conversely, de Wilde et al. froze Stable Diffusion and trained specific text prompts to generate 2D prostate MRI slices with or without suspicious lesions [17].
Synthetic medical image generation in 3D is faced with additional challenges related to the increased number of pixels per image compared to 2D LDMs. To combat LDM hallucinations in generating 3D CT images, Xu et al. split the synthetic image generation process into low- and high-resolution stages while generating both images and segmentation maps simultaneously in the low-resolution stage [18]. The upscaled segmentation maps then enabled higher-quality synthetic image generation in the high-resolution LDM stage. MAISI focused instead on computational efficiency and developed a secondary LDM to augment real segmentation masks for use in generating synthetic 3D CT images [1].
2.3 Image Conditioning with Diffusion Models
Conditioning allows for control of the denoising U-Net behavior during LDM training and inference. LDM conditioning typically takes up to three forms. The first is timestep embedding, informing the LDM of its stage in the denoising process. Similar to positional encoding in transformers [19], the timestep is encoded with sinusoidal encoding and then conditions the LDM through addition with the latent image. The second is prompt embedding, informing the LDM of additional information during denoising. Prompt embedding can take different forms but typically are images or text. Prompting data are encoded with separate encoders and condition the LDM through concatenation with the latent image (usually for image prompting) or by cross-attention at multiple layers within the denoising U-Net (usually for text prompting). Class conditioning is a lesser-used third method that expands upon the timestep embedding by adding or concatenating a vector with class information to the timestep embedding [10]. This allows for class-based information to condition the U-Net during denoising. For example, MAISI was trained to accept segmentation masks as prompt embedding, and included both image voxel spacing and body region as class conditioning that modified the timestep embedding [1].
LDM text conditioning typically relies on the Contrastive Language-Image Pre-Training (CLIP) text encoder [20], trained to encode text to a shared image-text latent space. Beyond being trained on non-medical text, the CLIP text encoder is limited by its 77 text token input restriction, with literature documenting an effective length of under 20 tokens [21]. Consequently, text prompting for commercial LDMs are typically restricted to keywords or short phrases. Within medicine however, clinical reports are typically longer and rich in free-text information including location, disease severity, or other defining characteristics that could improve the realism of synthetic image generation [6].
Strategies for text prompting in medical imaging LDMs include using short phrases [22], training new embeddings for the frozen CLIP text encoder [17], fine-tuning the CLIP text encoder on medical data [5], or training new LDMs with a domain-relevant medical LLM [5, 18].
Outside of the medical space, the CLIP text token restrictions have prompted innovation in adapting large language models (LLMs) for use with LDMs. Notably, the recent Efficient Large Language model Adapter (ELLA) demonstrated improvement in LDM performance in both image generation and long-prompt following [23]. ELLA received the output of the T5 XL LLM which was inputted to a chain of six Timestep-aware Semantic Connector (TSC) blocks to generate the final embedding to condition the LDM. Details on TSC block design can be found in [23]. The purpose of ELLA was to leverage the power of LLMs for text feature extraction without retraining either of the LLM or LDM.
PathLDM is the most similar in motivation to our proposed work [6]. Recognizing the richness of clinical text, the value in using class-based information, the utility of automated ground-truth labeling, and the limitations to the CLIP text encoder for conditioning LDMs, PathLDM developed an automated text prompt generation workflow for synthetic 2D breast cancer histopathology slide generation. This workflow extracted two pathology-related probabilities (“Tumor” and “TIL”) with pretrained models that were each assigned “high” or “low” after probability thresholding, and summarized each slide’s clinical report to under 75 words using the GPT-3.5 LLM [24]. PathLDM also replaced the CLIP text encoder with a pretrained, histopathology-specific CLIP text encoder variant [25]. The final text prompt for LDM conditioning consisted of two class-based keywords and the summarized report in the form “[high/low] Tumor, [high/low] TIL, [report summary]”, allowing PathLDM to receive both classifier and key text report information as conditioning.
Different from PathLDM, our method:
  • Uses text from the original report instead of a LLM-generated summary;
  • Trains the class extractor simultaneously with the LDM, rather than relying on a pretrained solution for class extraction;
  • Adapts a pretrained non-medical LLM for medical use instead of using a medical domain-specific CLIP text encoder; and
  • Performs class conditioning using timestep embedding rather than through text encoding.
Compared to ELLA, our method:
  • Modifies each TSC block by adding pathology class extractor layers;
  • Focuses additionally on class conditioning through the timestep embedding rather than solely on adapting the text embedding to the latent space;
  • Focuses on generating 3D medical images instead of 2D non-medical images; and
  • Trains the LDM U-Net from scratch alongside the adapter rather than using a pretrained LDM U-Net.
3 Methods
3.1 Dataset
This retrospective study used axial T2-weighted prostate MRI of men at high-risk for prostate cancer. This study was approved by an institutional research ethics board and informed consent was waived. The complete dataset used 4153 images from one healthcare institution plus 1658 images from two publicly available datasets (1500 from the PI-CAI Grand Challenge[13], 158 from the Prostate158 dataset[26]) for a total dataset size of 5811. Inclusion criteria for data from the healthcare institution center were referral for prostate MRI, being treatment naïve, and study date between April 2009 and August 2021. Exclusion criteria were incomplete or non-diagnostic image quality. DICOM images from the healthcare institution were converted to NIFTI format with patient information removed from image metadata prior to this study. Radiology report texts were preprocessed to remove identifying information including patient name or doctor initials. Data inclusion and exclusion criteria for each public dataset have been published[13, 26]. The public datasets were already anonymized.
A hold-out test set of 740 images was randomly selected on a per-case level and used to assess the results in this study. Clinical reports were available for 4150 studies from the healthcare institution. No public dataset used in this study included full clinical reports with their images. Because some images from the healthcare institution were acquired prior to the latest update in prostate MRI reporting standard, these cases were reviewed and assigned a new pathology classification per the updated reporting guidelines [27]. All patients had either a radiology classification or histopathology results from within 6 months of the MRI. In the event a radiology classification was not present for an image within the training set, the histopathology results were used instead. Images were center-cropped to a $128mm\times 128mm$ field of view in the axial plane, normalized and then resampled to $256\times 256\times 128$ voxel dimensions using Lanczos interpolation [28].
3.2 Implementation
The proposed framework in this study uses two pretrained components. The first is the MAISI LDM VQ-VAE, pretrained on 3D CT and MRI images [1]. The second is Google’s FLAN-T5 XXL LLM [11]. The base LDM in this study used the frozen MAISI VQ-VAE with a learnable 4-channel, 4-layer denoising U-Net that concatenated image voxel spacings to the timestep embedding. The frozen Flan-T5 XXL text encoder was used to encode the clinical report text.
CCELLA extends the ELLA adapter from [23] by adapting the encoded text for both deep cross-attention and timestep embedding conditioning. CCELLA begins with six class-conditioning TSC (CC-TSC) blocks, each consisting of a TSC block with a separate classifier head of two fully-connected layers of hidden size 256 and output size 2 (MLP block) at the end of the block. The outputs from all CC-TSC classifier heads were concatenated and fed into a final fully-connected layer to generate an output of size 2 to be concatenated with the timestep embedding (Figure 1).
Figure 1: Overview of CCELLA. Consists of six CC-TSC blocks with classifier and aligned text embedding heads. TSC block description can be found in [9].
To train CCELLA with the denoising U-Net, a new joint loss function is proposed:
$L_{CCELLA}=L_{Reconstruction}+\lambda L_{Class}$ (1)
The intent of this loss function is to emphasize both LDM image reconstruction as well as appropriate pathology extraction from the reference text. For this study, we used L1 loss for reconstruction and Focal Loss for class prediction, with class loss weighting $\lambda=10^{-4}$.
3.3 Training and Assessment
CCELLA design can be found in Figure 1, providing both cross-attention and timestep conditioning to the LDM. CCELLA performance was assessed against three baseline LDMs: 1) MAISI without text or class conditioning (BASE), 2) MAISI with radiology classification timestep conditioning (RADBASE), 3) MAISI with text conditioning using FLAN-T5 XXL and a chain of six TSC blocks (ELLABASE). All LDMs were trained for 400 epochs with initial learning rate of $10^{-4}$, the AdamW optimizer, a second-order polynomial learning rate scheduler, and batch size of 16 on four Tesla V100 GPUs. All models received image voxel spacing as timestep conditioning as per the foundation MAISI model [1]. During the training of LDMs with text conditioning, images that did not have a text report used an empty string as a proxy text report. When training CCELLA, $\lambda$ in (1) was set to 0 for these cases so that the classifier portion of the loss function did not update the model for these images.
Fréchet Inception Distance (FID) was used to evaluate LDM performances [29]. FID was calculated in two ways: 1) in 3-dimensions using the Med3D ResNet50 pretrained on 23 datasets [30], 2) in 2-dimensions for each of the sagittal, coronal and axial planes using a ResNet50 pretrained on RadImageNet [31].
3.4 Ablation Studies
Two ablation studies were performed. The first assessed the components of CCELLA, namely the joint conditioning and classifier heads and the joint loss function. This ablation study compared CCELLA to 1) an LDM with ELLA text conditioning and separate timestep conditioning using ground truth radiology classification (DECOUPLED), 2) CCELLA using only L1 loss instead of the modified loss in (1) (L1ONLY). The second ablation study compared CCELLA to variants by: 1) placing only one CC-TSC block after a chain of five TSC blocks (LAST), 2) placing one CC-TSC block before five TSC blocks (FIRST), 3) chaining six CC-TSC blocks where class adapter weights are shared across all blocks (SHARED).
3.5 Downstream Classification Task
The trained CCELLA LDM generated one synthetic image for every real training sample using the sample’s clinical text. Both the radiologist label from the source training sample (“clinical label”) and the CCELLA-generated label (“synthetic label”) were stored for each synthetic image during generation. An EfficientNet-b0 [32] was trained in each of five scenarios: 1) real data, 2) synthetic data with clinical labels, 3) synthetic data with synthetic labels, 4) real and synthetic data with clinical labels, 5) real and synthetic data with synthetic labels for synthetic images. For each scenario, the same real and synthetic data were used in training (i.e. synthetic data was not regenerated between scenarios). Models were trained across four Tesla V100 GPUs with a batch size of 16 for 16k steps at an initial learning rate of $10^{-4}$ with the AdamW optimizer and a second-order polynomial learning rate scheduler. Trained EfficientNet-b0 performances were assessed using classification accuracy on the same unseen test set as the LDM evaluation. Secondary assessment was performed using sensitivity, specificity, positive predictive value (PPV), and negative predictive value (NPV).
4 Results
4.1 Performance Comparison
Table 1: CCELLA Performance vs Baselines
Model 3D FID $\downarrow$ 2D Sagittal FID $\downarrow$ 2D Coronal FID $\downarrow$ 2D Axial FID $\downarrow$
BASE 0.071 10.03 8.75 13.24
RADBASE 0.077 10.96 9.53 14.10
ELLABASE 0.080 11.64 10.40 15.20
DECOUPLED 0.048 8.33 7.58 11.57
L1ONLY 0.051 7.93 7.21 10.94
CCELLA (ours) 0.025 5.27 4.73 8.09
Table 1 shows the FID performance of CCELLA against its baseline and ablation models. CCELLA outperformed all other models in both 3D and all-axis 2D FID metrics. Providing no conditioning, or only one of timestep or cross-attention conditioning performed worse than providing both. Example images generated by CCELLA can be seen in Figure 2.
Figure 2: Four sample axial T2 prostate MRI images generated by CCELLA
4.2 Ablation Study
Table 2: Ablation Study: CCELLA Variants
Model 3D FID $\downarrow$ 2D Sagittal FID $\downarrow$ 2D Coronal FID $\downarrow$ 2D Axial FID $\downarrow$ Accuracy $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$ PPV $\uparrow$ NPV $\uparrow$
SHARED 0.065 7.84 7.05 11.01 0.83 0.82 0.85 0.87 0.79
FIRST 0.053 10.60 9.33 14.16 0.82 0.79 0.86 0.88 0.76
LAST 0.051 8.30 7.16 11.34 0.82 0.79 0.85 0.87 0.77
CCELLA 0.025 5.27 4.73 8.09 0.81 0.77 0.86 0.87 0.75
Table 2 shows the FID and confusion matrix performances of CCELLA against its variants. CCELLA outperformed its variants in all FID metrics. Interestingly, CCELLA performed worse than its variants at class prediction from text, with the SHARED variant achieving best class prediction accuracy.
4.3 Downstream Classification Task
Table 3: Classification Results when Trained with Combinations of Real Data (R), Synthetic Data with Clinical Labels ($S_{R}$), and Synthetic Data with Synthetic Labels ($S_{S}$)
Model AUC AP Accuracy Sensitivity Specificity PPV NPV % Pos
$R$ 0.76 0.73 0.69 0.67 0.72 0.75 0.63 48%
$S_{R}$ 0.77 0.77 0.64 0.51 0.81 0.77 0.57 48%
$S_{S}$ 0.80 0.80 0.66 0.50 0.87 0.83 0.58 39%
$R+S_{R}$ 0.80 0.79 0.74 0.71 0.77 0.80 0.68 48%
$R+S_{S}$ 0.79 0.79 0.71 0.71 0.72 0.76 0.66 43%
Table 3 shows the performance of EfficientNet-b0 classifiers trained for identifying suspicious lesions in prostate MRI trained on any of real ($R$), synthetic with clinical label ($S_{R}$), and synthetic with synthetic label ($S_{S}$) datasets. Area under the receiver operating characteristic curve (AUC) and average precision (AP) report diagnostic performance agnostic of prediction threshold. Accuracy (Acc.), sensitivity (Sens.), specificity (Spec.), positive predictive value (PPV) and negative predictive value (NPV) were calculated at a threshold of 0.5. Augmenting real data with synthetic data for training improved all classifier metrics regardless of synthetic ground truth type, with the exception of specificity when using synthetic data with synthetic labels which remained the same. Training on synthetic data using synthetic labels as ground truth alone achieved highest AUC, AP, specificity and PPV among classifiers assessed. CCELLA generated a negative classification by default when text was missing from an image, resulting in a lower proportion of positive labels (Pos) for training for $S_{S}$. All 468 cases in the training dataset where CCELLA extracted negative classification from a positive source image occurred when clinical report text was missing. The test set had a positive case percentage of 56%.
5 Discussion
This study sought to introduce a new framework for training higher-quality synthetic prostate MRI using a limited and fixed volume dataset with limited supervised labeling. This study achieved this by proposing CCELLA and a joint loss function, in addition to reusing pretrained components where possible. Our proposed strategy achieved higher performance than baseline LDMs conditioned on nothing, on radiology label only, and on text only using ELLA. The first ablation study sought to understand the mechanism by which CCELLA achieved better performance. This ablation study trained: 1) an LDM receiving text conditioning by cross-attention using ELLA and timestep conditioning using the ground truth radiology label, 2) an LDM fitted with CCELLA but without using the proposed joint loss. CCELLA outperformed both of these variants suggesting that it is not simply the addition of a class label in the timestep embedding (DECOUPLED) nor the extension of the text adapter for timestep conditioning (L1ONLY) that results in CCELLA’s success. These results suggest that the joint training of the classifier with the text adapter encourages CCELLA to extract more clinically relevant information from text encoded by a non-clinically-trained LLM. While further investigation should be performed to confirm these findings, this ablation study suggests that pretrained LLMs outside of the medical space could be better adapted for use on medical data when trained jointly with clinical labels in lieu of adapting a pretrained, non-medical LLM without additional labels or training a new LLM on clinical information.
CCELLA was also compared against alternate designs in a second ablation study. Interestingly, these results indicate a near-inverse relationship between diffusion output quality and classifier accuracy, with CCELLA achieving the best in output quality but the worst in classifier accuracy and the SHARED variant achieving the worst in image generation but the best in classifier accuracy. While an initial intuition of these results might suggest that there is a trade-off between the classifier and text adapter performances, further investigation should seek to further explore this relationship and understand how it might be better leveraged for both improved LDM and classifier performance.
Sample images from the CCELLA-trained LDM can be seen in Figure 2. Qualitatively, these images exhibit natural variance in the sizes and shapes of the prostate while consistently depicting the different anatomical regions within the prostate. Some images display differences in brightness or smoothness of the image which can give a slightly less “realistic” appearance despite the anatomical accuracy and the utility of these images in augmenting the dataset of a downstream task. Thus, there exist possible research opportunities in developing strategies for the measuring and improving the consistency of image quality during synthetic medical image generation with diffusion models.
Finally, this study trained downstream classifier models for prostate cancer prediction from MRI using real data and synthetic data generated by a CCELLA-trained LDM. The $R+S_{R}$ and $R+S_{S}$ classifiers performed equally or better than the $R$ classifier on all metrics, with $R+S_{R}$ outperforming $R$ on all confusion matrix metrics. These results validate the hypothesis that synthetic data can be used to successfully augment the training dataset for a classifier model and aligns with recent literature that generate ground truths alongside data samples. $S_{S}$ also performed comparably to $R$ with higher AUC and AP but lower accuracy, suggesting that the CCELLA-generated synthetic data may be as useful as real data when training downstream machine learning tasks after tuning the classifier threshold. These results are promising, though it is important to note that the synthetic data was generated from the same distribution as the real data. Future work should seek to investigate these results further by generating synthetic data for training with real data of different distribution.
Interestingly and in contrast to expectation, the $S_{S}$ classifier, trained only on the same synthetic data as $S_{R}$ but with ground truth labeling sourced from CCELLA’s classifier head instead of the clinical label accompanying the source image, outperformed not only the $S_{R}$ classifier in all metrics except sensitivity, but also the $R$ classifier in multiple metrics (AUC, AP, specificity, PPV). Seemingly paradoxically, augmenting the real dataset with $S_{R}$ resulted in better performance than augmenting with $S_{S}$ in all metrics except AP and sensitivity where they were equal. This likely occurred for three reasons. Firstly, only 3701 of the 5071 training samples had clinical report text. Because null text reports generated negative class predictions by default, the label distribution for synthetic labeling was different from from the real label distribution which would have affected the model training. It therefore makes retrospective sense that $R+S_{S}$, having trained on a dataset with lower positive case percentage than $R+S_{R}$ (43% versus 48%), would exhibit more difficulty in adapting to a test set with higher positive case percentage (56%), resulting in lower accuracy (0.71 versus 0.74). Secondly and related to the first, the training cases that were missing radiology text belonged to both positive and negative classes. Thus even though the CCELLA classifier head was not trained on these cases, the LDM would still have learned to generate both positive and negative images when presented with a null prompt. This in-turn could have affected the performance of classifiers trained with synthetic data. Thirdly, the CCELLA LDM did not receive the real image as input during inference. Consequently, even though the ground-truth radiology classification was based on the radiologist re-interpretation of the full image under the most recent prostate MRI reporting standard instead of simply extracting the classification in the clinical text, the use of the ground truth re-classification may have lost its meaning in favor of what the LLM interprets from the clinical text. This could have resulted in the synthetic images aligning well with the text interpretation and CCELLA class extraction but not necessarily with the radiologist re-classification of the real image corresponding to that text, which could explain why $S_{R}$ exhibited poorer performance. Thus, while the same train and test datasets used for LDM training were used for the downstream classifier task for the sake of consistency, future work should seek to further explore the difference between using real versus synthetic labels for synthetic images for the purpose of dataset augmentation when training a model for a medical task on data.
This study has a few limitations, each of which present an opportunity for future work. Firstly, this study emphasized data from a single institution, for a single modality in a single pathology. While the pathology and institution restrictions were intentional to the study design in developing a limited-data LDM training framework, future work should seek to adapt this framework to multi-institutional data, as well as images of different modality or pathology to assess method generalizability. Additionally, prostate MRI is clinically assessed in a minimum of three, and preferably four different MRI sequences [33, 34]. Consequently, portions of the clinical text used for conditioning would have referred to MRI sequences that the LDM was neither trained on nor asked to generate. While the use of one sequence alone in this study was chosen to focus primarily on the adapter innovation and data-limited framework design instead of the additional challenge of multi-sequence generation, future work should seek to extend data-limited LDM training strategies for optimal multi-sequence generation. As well, this study included medical images from public datasets that did not have accompanying text reports for the purpose of gathering more training samples to train the denoising U-Net and the downstream classifier models. While outside the scope of this study, this prompts two opportunities for future work. One opportunity is to investigate how the ratio of conditioned to unconditioned data samples affect model training and whether there is a critical ratio below which training with conditioning becomes unstable or negatively affects training. Another opportunity is to explore synthetic clinical text report generation for these data samples, whereby either a rule-based algorithm or LLM informed with the radiologist classification and image metadata provided in the public dataset would generate text reports for use in conditioning the LDM during training and inference. Finally, while the intent of this study aligns with a growing body of literature seeking to use LDMs to generate synthetic medical data for the purpose of mitigating data scarcity challenges within machine learning for medical imaging applications, recent literature has also shown that LDMs can occasionally memorize and reproduce their training samples on inference [35]. While outside the scope of this study, future work investigating and improving upon differential privacy within medical imaging LDMs would be valuable.
6 Conclusion
In summary, this study presents CCELLA, a novel adapter for simultaneous alignment of non-medical LLMs to the medical image space and pathology extraction from medical text for LDM conditioning, as well as a CCELLA-centric framework for medical image LDM training given limited data volume and annotation. Our proposed method outperforms comparative baseline models when generating synthetic prostate MRI data and successfully augments the training dataset of a downstream classifier model for improved cancer prediction performance over training on real data alone. The results of this study suggest that CCELLA can successfully adapt pretrained, non-medical LLMs for use in conditioning LDMs for medical image generation, and can successfully train a 3D medical image LDM using data generated during clinical routine. Future work should focus on expanding this framework to multi-sequence image generation, validating the framework in a different medical imaging and pathology space, and exploring differential privacy within medical imaging LDMs.