### v2

# Prompt-Guided Latent Diffusion with Predictive Class Conditioning for 3D Prostate MRI Generation

Abstract
Objective: Latent diffusion models (LDM) could alleviate data scarcity challenges affecting machine learning development for medical imaging. However, medical LDM strategies typically rely on short-prompt text encoders, non-medical LDMs, or large data volumes. These strategies can limit performance and scientific accessibility. We propose a novel LDM conditioning approach to address these limitations. Methods: We propose Class-Conditioned Efficient Large Language model Adapter (CCELLA), a novel dual-head conditioning approach that simultaneously conditions the LDM U-Net with free-text clinical reports and radiology classification. We also propose a data-efficient LDM framework centered around CCELLA and a proposed joint loss function. We first evaluate our method on 3D prostate MRI against state-of-the-art. We then augment a downstream classifier model training dataset with synthetic images from our method. Results: Our method achieves a 3D FID score of 0.025 on a size-limited 3D prostate MRI dataset, significantly outperforming a recent foundation model with FID 0.071. When training a classifier for prostate cancer prediction, adding synthetic images generated by our method during training improves classifier accuracy from 69% to 74%. Training a classifier solely on our method’s synthetic images achieved comparable performance to training on real images alone. Conclusion: We show that our method improved both synthetic image quality and downstream classifier performance using limited data and minimal human annotation. Significance: The proposed CCELLA-centric framework enables radiology report and class-conditioned LDM training for high-quality medical image synthesis given limited data volume and human data annotation, improving LDM performance and scientific accessibility. Code from this study will be available at https://github.com/grabkeem/CCELLA
Keywords Generative AI  $\cdot$ Large Language Model Adapters  $\cdot$ Medical Image Generation  $\cdot$ Magnetic Resonance Imaging  $\cdot$ Prostate Cancer
1 Introduction
Machine learning (ML) for medical image interpretation is a rapidly expanding field intent on improving imaging-based healthcare. However, recent ML advancements outside of medical imaging require training data in higher orders of magnitude than their predecessors. These advances cannot translate easily to the medical imaging space due to labeled data scarcity challenges including patient privacy and the cost of expert annotation. Recent medical imaging literature has seen an increase in latent diffusion model (LDM) studies [1]. Artificially increasing ML training volume with synthetic data produced by these generative methods provides a possible solution to labeled training data scarcity [2].
Developing LDMs for medical imaging poses a few challenges. Training LDMs from scratch require large volumes of data. The Stable Diffusion LDM [3] for generating nonmedical, 2D color images was trained initially on 400 million image-text pairs [4]. A subsequent version of Stable Diffusion trained on 5.85 billion image-text pairs [5]. These data volumes are inaccessible to most healthcare institutions due to aforementioned data scarcity challenges. Non-medical 2D LDMs have been successfully fine-tuned on 2D medical images [6], however this strategy cannot easily extend to 3D. Adding labeled data conditioning has also shown to increase data efficiency during LDM training [7]. To this end, some medical LDMs have used the pre-trained CLIP text encoder to condition the LDM, similar to Stable Diffusion [3]. However, research has shown that the CLIP text encoder mostly emphasizes the first 20 tokens [8]. Thus, these medical LDM studies typically use only short, curated medical phrases or keywords with either the existing [9] or fine-tuned [6] CLIP text encoder rather than using the detail-rich clinical reports. These strategies require extra annotation and preprocessing or volumes of data that are challenging for a single institution to gather. This consequently limits the scientific accessibility of medical LDM development.
In this study we propose Class-Conditioned Efficient Large Language Model Adapter (CCELLA), a text adapter extending ELLA [10], to improve LDM performance for synthetic medical image generation in a challenging medical imaging domain. CCELLA processes the text embedding generated by a large language model (LLM) from clinical reports prior to dual-head conditioning of the LDM U-Net. CCELLA performs two functions simultaneously: the first head aligns the text embedding with the image latent space prior to conditioning; the second head provides radiologist-derived class conditioning for the LDM during training. Each of these functions have independently demonstrated improvements in LDM performance outside of medical imaging [10, 11]. This joint function is hypothesized to provide two advantages: 1) to assert that the LDM receives conditioning about the presence or absence of suspicious lesions within the image, 2) more importantly, to inform CCELLA which text features should be emphasized when aligning the embedding to the image latent space for LDM conditioning. Notably, it is hypothesized that performing both functions in a single adapter will improve LDM performance beyond simply performing both functions independently. To achieve these functions, we also propose a modified loss function. This loss function simultaneously trains the LDM to produce high quality medical images and to appropriately extract radiology classification information from the clinical report text.
Finally, we propose an LDM training framework around CCELLA that uses limited human data annotation and a training data volume reasonably accessible to a single healthcare institution. The intent of this framework is to improve the scientific accessibility of LDMs for medical imaging despite limited available data volumes. The framework functions by exploiting the clinical report and labeling already performed by radiologists during clinical routine and by reducing training requirements with publicly available, pre-trained components [2, 12].
This study focused specifically on MRI for prostate cancer as the target domain to evaluate this framework. Prostate MRI is scored by radiologists using the Prostate Imaging – Reporting and Data System (PI-RADS) guidelines [13]. This pathology and imaging method was selected due to its difficulty in clinical interpretation, its moderate inter-reader variability between clinical experts, and the relatively low-availability of publicly available domain-specific data with the largest dataset consisting of 1500 samples [14, 15]. This study then performed a series of ablations to assess the mechanism by which CCELLA improves LDM performance and compared performance differences among variants of the proposed adapter. To validate utility of the proposed LDM training framework for dataset augmentation, this study demonstrated that the synthetic data generated with our proposed LDM framework can improve a classifier model at interpreting prostate MRI for cancer presence.
In summary, this paper proposes the following:
  • A novel dual-head adapter for simultaneous alignment of medical text features extracted by a non-medical LLM to the medical imaging space and PI-RADS score extraction from medical text
  • A LDM training framework for improved radiology class-conditioned synthetic medical image generation that is trainable with a reasonable data volume.
  • A method for augmenting datasets to improve downstream ML tasks using LDM-generated synthetic images with limited data annotation beyond clinical routine.
2 Related Work
2.1 Latent Diffusion Models
Latent diffusion models (LDM) are a type of generative model consisting of two stages: a vector-quantized variational autoencoder (VQ-VAE) and a denoising U-Net [3]. The VQ-VAE compresses the image to a lower-dimensional representation in a discretized latent space [16]. Functionally this reduces computational complexity increasing both training and inference speed, and coerces the image representation to fit a normal distribution. The denoising U-Net learns the reverse process to a Markov Chain [17]. During LDM training, the Markov Chain forward process gradually adds Gaussian noise to the latent image for each of many, e.g. 1000, timesteps. The denoising U-Net is then trained to reverse the noise addition process by predicting the noise added between the true latent image and a randomly selected timestep up to 1000.
At inference time, the synthetic latent image begins as randomly-generated Gaussian noise. The U-Net then proceeds to denoise the image from timestep 1000 to 0, at which point the resultant latent image is decoded by the VQ-VAE to yield a synthetic image.
2.2 Conditioning and Medical Diffusion Models
LDMs have shown promising results in generating synthetic images both outside and within medical imaging [1]. Medical imaging LDM studies typically require either large data volumes to train from scratch, or strategic data-efficient methods to adapt pretrained LDMs. One critical, strategic approach is to condition the LDM with additional information (e.g. segmentation maps, text, or PI-RADS class) for improved image generation performance [18, 19, 2].
Conditioning allows for control of the denoising U-Net behavior during LDM training and inference. LDM conditioning typically takes up to three forms. The first is timestep embedding, informing the LDM of its stage in the denoising process. Similar to positional encoding in transformers [20], the timestep is encoded with sinusoidal encoding and then conditions the LDM through addition with the latent image. The second is prompt embedding, informing the LDM of additional information during denoising. Prompt embedding can take different forms but typically are images or text. Prompting data are encoded with separate encoders and condition the LDM through concatenation with the latent image (usually for image prompting) or by cross-attention at multiple layers within the denoising U-Net (usually for text prompting). Class conditioning is a lesser-used third method that expands upon the timestep embedding by adding or concatenating a vector with class information to the timestep embedding [11]. This allows for class-based information to condition the U-Net during denoising. These conditioning methods can occur in parallel. For example, MAISI was trained to accept segmentation masks as prompt embedding, and included both image voxel spacing and body region as class conditioning that modified the timestep embedding [2].
LDM text conditioning typically relies on the Contrastive Language-Image Pre-Training (CLIP) text encoder [21], trained to encode text to a shared image-text latent space. The CLIP text encoder is limited by its 77 text token input restriction, with literature documenting an effective length of under 20 tokens [22]. Consequently, text prompting for commercial LDMs are typically restricted to keywords or short phrases. This encoder was also not trained specifically for use with medical text. By contrast, clinical reports are typically longer and rich in free-text information (e.g. location, disease severity, or other defining characteristics) that could improve the realism of synthetic image generation [7]. Consequently, literature has developed multiple strategies for text prompting in medical imaging LDMs. These include crafting short phrases [23], training new embeddings for the frozen CLIP text encoder [18], fine-tuning the CLIP text encoder on medical data [6], or training new LDMs with a domain-relevant medical LLM [6, 19].
Outside of the medical space, the CLIP text token restrictions have prompted innovation in adapting large language models (LLMs) for use with LDMs. Notably, the recent Efficient Large Language model Adapter (ELLA) demonstrated improvement in LDM performance in both image generation and long-prompt following [24]. ELLA received the output of the T5 XL LLM which was inputted to a chain of six Timestep-aware Semantic Connector (TSC) blocks to generate the final embedding to condition the LDM. Details on TSC block design can be found in [24]. The purpose of ELLA was to leverage the power of LLMs for text feature extraction without retraining either of the LLM or LDM.
PathLDM [7] is the most similar in motivation to our proposed work, as the authors of that study recognized the richness of clinical text, the value in using class-based information, the utility of automated ground-truth labeling, and the limitations to the CLIP text encoder for conditioning LDMs. PathLDM developed an automated text prompt generation workflow for synthetic 2D breast cancer histopathology slide generation. This workflow extracted two pathology-related probabilities (“Tumor” and “TIL”) with pretrained models that were each assigned “high” or “low” after probability thresholding, and summarized each slide’s clinical report to under 75 words using the GPT-3.5 LLM [25]. PathLDM also replaced the CLIP text encoder with a pretrained, histopathology-specific CLIP text encoder variant [26]. The final text prompt for LDM conditioning consisted of two class-based keywords and the summarized report in the form “[high/low] Tumor, [high/low] TIL, [report summary]”, allowing PathLDM to receive both classifier and key text report information as conditioning.
Different from PathLDM, our method:
  • Uses text from the original report instead of a shorter, LLM-generated summary;
  • Trains the class extractor simultaneously with the LDM, rather than relying on a non-medical, pretrained solution;
  • Adapts a pretrained non-medical LLM for medical use instead of using a medical domain-specific CLIP text encoder; and
  • Performs class conditioning using timestep embedding rather than through text encoding to enable class-specific training of the text adapter.
Compared to ELLA, our method:
  • Modifies each TSC block by adding PI-RADS score extractor layers;
  • Focuses additionally on class conditioning through the timestep embedding rather than solely on adapting the text embedding to the latent space;
  • Focuses on generating 3D medical images instead of 2D non-medical images; and
  • Trains the LDM U-Net from scratch alongside the adapter rather than using a pretrained LDM U-Net.
3 Methods
3.1 Dataset
This retrospective study used axial T2-weighted prostate MRI of men at high-risk for prostate cancer. This study was approved by an institutional research ethics board and informed consent was waived. The complete dataset used 4153 images from one healthcare institution plus 1658 images from two publicly available datasets (1500 from the PI-CAI Grand Challenge [15], 158 from the Prostate158 dataset [27]) for a total dataset size of 5811. Inclusion criteria for data from the healthcare institution center were referral for prostate MRI, being treatment naïve, and study date between April 2009 and August 2021. Exclusion criteria were incomplete or non-diagnostic image quality. DICOM images from the healthcare institution were converted to NIFTI format with patient information removed from image metadata prior to this study. Radiology report texts were preprocessed to remove identifying information including patient name or doctor initials. Data inclusion and exclusion criteria for each public dataset have been published [15, 27]. The public datasets were already anonymized.
[…]
3.2 Latent Diffusion Model Implementation
The proposed LDM framework uses two pretrained components. The first is the MAISI LDM VQ-VAE, pretrained on 3D CT and MRI images [2]. The second is the FLAN-T5 XXL LLM [12]. The base LDM in this study used the frozen MAISI VQ-VAE with a learnable 4-channel, 4-layer denoising U-Net that concatenated image voxel spacings to the timestep embedding. The frozen FLAN-T5 XXL text encoder was used to encode the clinical report text.
CCELLA extends the ELLA adapter from [24] by adapting the encoded text for both deep cross-attention and timestep embedding conditioning. CCELLA begins with six class-conditioning TSC (CC-TSC) blocks, each consisting of a TSC block with a separate classifier head of two fully-connected layers of hidden size 256 and output size 2 (MLP block) at the end of the block. The outputs from all CC-TSC classifier heads were concatenated and fed into a final fully-connected layer to generate an output of size 2 to be concatenated with the timestep embedding. The image voxel spacing was additionally concatenated to the timestep embedding as per the foundation MAISI model [2]. The output of the final CC-TSC text embedding head conditioned the denoising U-Net through cross-attention. Figure 1 illustrates the CCELLA adapter design as well as the full dataflow for the LDM framework. Class conditioning was intentionally implemented as a separate vector for timestep embedding instead of through additional text prompt words as performed in PathLDM [7]. Our implementation enabled direct training of the CCELLA adapter using radiologist classification rather than relying on the adapter to train implicitly for improved image generation alone.
Figure 1: Overview of CCELLA-centric framework (left) and CCELLA adapter (right). Image encoder ($E_{Image}$) and decoder ($D_{Image}$) from [2]. Text encoder ($E_{Text}$) from [12]. CCELLA consists of six CC-TSC blocks with classifier ($h_{C}$) and aligned text embedding ($h_{T}$) heads. TSC block originally from [10]. Framework input includes the radiology report text for training and inference, with image and radiology classification additionally included for training alone. Example text redacted for privacy.
To train CCELLA with the denoising U-Net, a new joint loss function is proposed:
$L_{CCELLA}=L_{Reconstruction}+\lambda L_{Class}$ (1)
The intent of this loss function is to emphasize both LDM image reconstruction as well as appropriate PI-RADS score extraction from the reference text. For this study, we used L1 loss for reconstruction and focal loss for class prediction, with class loss weighting $\lambda=10^{-4}$. L1 loss was selected due to its frequent use in LDM image reconstruction [3, 2]. Focal loss was selected for its success in the PI-CAI challenge [15]. Class loss weighting was determined empirically after training on a subsample of the full training set.
3.3 CCELLA Training and Evaluation
The MAISI and ELLA studies have already demonstrated success against state-of-the-art models [2, 24]. Specifically, the MAISI framework demonstrated improved Fréchet Inception Distance (FID) in all 2-dimensional planes compared to a pretrained, medical generative adversarial network [2]. ELLA demonstrated qualitative and quantitative improvement when adding their adapter to an existing commercial LDM [24]. Thus for this study, we assessed the CCELLA-centric framework against: 1) the MAISI LDM (MAISI) without text or class conditioning, 2) MAISI with radiology classification timestep conditioning (RADMAISI), 3) MAISI with text conditioning using ELLA to adapt FLAN-T5 XXL (ELLAMAISI).
Two ablation studies were then performed. The first assessed the components of CCELLA, namely the text and classifier conditioning heads as well as the joint loss function. This ablation study compared CCELLA to 1) an LDM with ELLA text conditioning and separate timestep conditioning using ground truth radiology classification (DECOUPLED), 2) CCELLA using only L1 loss instead of the proposed joint loss in (1) (L1ONLY). The second ablation study compared the CCELLA design to variants by: 1) placing only one CC-TSC block after a chain of five TSC blocks (LAST), 2) placing one CC-TSC block before five TSC blocks (FIRST), 3) chaining six CC-TSC blocks where class adapter weights were shared across all blocks (SHARED).
All models received image voxel spacing as timestep conditioning as per the foundation MAISI model [2].
Training details: All images were center-cropped to a $128mm\times 128mm$ field of view in the axial plane, normalized and then resampled to $256\times 256\times 128$ voxel dimensions using Lanczos interpolation [28] prior to LDM training and evaluation. All LDMs were trained for 400 epochs with initial learning rate of $10^{-4}$, the AdamW optimizer, a second-order polynomial learning rate scheduler, and batch size of 16 on four Tesla V100 GPUs. During the training of LDMs with text conditioning, images that did not have a text report used an empty string as a proxy text report. When training CCELLA, $\lambda$ in (1) was set to 0 for these cases so that the classifier portion of the loss function did not update the model for these images.
Evaluation: FID was used to evaluate LDM performances [29]. FID was calculated in two ways: 1) in 3-dimensions using the Med3D ResNet50 pretrained on 23 datasets [30], 2) in 2-dimensions for each of the sagittal, coronal and axial planes using a ResNet50 pretrained on RadImageNet [31].
3.4 Downstream Classification Task
The trained CCELLA LDM generated one synthetic image for every real training sample using the sample’s clinical text. Both the radiologist label from the source training sample (“clinical label”) and the CCELLA-generated label (“synthetic label”) were stored for each synthetic image during generation. An EfficientNet-b0 [32] was trained in each of five scenarios: 1) real data, 2) synthetic data with clinical labels, 3) synthetic data with synthetic labels, 4) real and synthetic data with clinical labels, 5) real and synthetic data with synthetic labels for synthetic images. For each scenario, the same real and synthetic data were used in training (i.e. synthetic data was not regenerated between scenarios).
Training details: All synthetic images were resampled to the original voxel size of their corresponding real image using Lanczos interpolation [28]. All images were linearly rescaled from 0%-98% intensity with clipping then padded to a $128mm\times 128mm\times 96mm$ field of view. Images were augmented during training with random flips, scaling, translations, and rotations. All classifiers were trained on four Tesla V100 GPUs with a batch size of 16 for 16k steps at an initial learning rate of $10^{-4}$ with the AdamW optimizer and a second-order polynomial learning rate scheduler.
Evaluation: Trained EfficientNet-b0 performances were assessed using the area under the receiver operating characteristic curve (AUC), average precision (AP), and classification accuracy on the same unseen test set as the LDM evaluation. Secondary assessment was performed using sensitivity, specificity, positive predictive value (PPV), and negative predictive value (NPV).
4 Results
4.1 Performance Comparison
Table 1: CCELLA Performance vs Baselines
Model 3D FID $\downarrow$ 2D Sagittal FID $\downarrow$ 2D Coronal FID $\downarrow$ 2D Axial FID $\downarrow$
MAISI 0.071 10.03 8.75 13.24
RADMAISI 0.077 10.96 9.53 14.10
ELLAMAISI 0.080 11.64 10.40 15.20
DECOUPLED 0.048 8.33 7.58 11.57
L1ONLY 0.051 7.93 7.21 10.94
CCELLA (ours) 0.025 5.27 4.73 8.09
Table 1 shows the FID performance of CCELLA against its baseline and ablation models. CCELLA outperformed all other models in both 3D and all-axis 2D FID metrics. Providing no conditioning, or only one of timestep or cross-attention conditioning performed worse than providing both. Example images generated by CCELLA can be seen in Figure 2.
Figure 2: Four real axial T2 prostate MRI images and the synthetic MRI images generated by ELLAMAISI and CCELLA conditioned on the same corresponding report text.
4.2 Ablation Study
Table 2: Ablation Study: CCELLA Variants
Model 3D FID $\downarrow$ 2D Sagittal FID $\downarrow$ 2D Coronal FID $\downarrow$ 2D Axial FID $\downarrow$ Accuracy $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$ PPV $\uparrow$ NPV $\uparrow$
SHARED 0.065 7.84 7.05 11.01 0.83 0.82 0.85 0.87 0.79
FIRST 0.053 10.60 9.33 14.16 0.82 0.79 0.86 0.88 0.76
LAST 0.051 8.30 7.16 11.34 0.82 0.79 0.85 0.87 0.77
CCELLA 0.025 5.27 4.73 8.09 0.81 0.77 0.86 0.87 0.75
Table 2 shows the FID and confusion matrix performances of CCELLA against its variants. CCELLA outperformed its variants in all FID metrics. Interestingly, CCELLA performed worse than its variants at class prediction from text, with the SHARED variant achieving best class prediction accuracy.
4.3 Downstream Classification Task
Table 3: Classification Results when Trained with Combinations of Real Data (R), Synthetic Data with Clinical Labels ($S_{R}$), and Synthetic Data with Synthetic Labels ($S_{S}$)
Model AUC AP Acc. Sens. Spec. PPV NPV % Pos
$R$ 0.76 0.73 0.69 0.67 0.72 0.75 0.63 48%
$S_{R}$ 0.77 0.77 0.64 0.51 0.81 0.77 0.57 48%
$S_{S}$ 0.80 0.80 0.66 0.50 0.87 0.83 0.58 39%
$R+S_{R}$ 0.80 0.79 0.74 0.71 0.77 0.80 0.68 48%
$R+S_{S}$ 0.79 0.79 0.71 0.71 0.72 0.76 0.66 43%
Table 3 shows the performance of EfficientNet-b0 classifiers trained for identifying suspicious lesions in prostate MRI trained on any of real ($R$), synthetic with clinical label ($S_{R}$), and synthetic with synthetic label ($S_{S}$) datasets. AUC and AP report diagnostic performance agnostic of prediction threshold. Accuracy (Acc.), sensitivity (Sens.), specificity (Spec.), positive predictive value (PPV) and negative predictive value (NPV) were calculated at a decision threshold of 0.5. Augmenting real data with synthetic data for training improved all classifier metrics regardless of synthetic ground truth type, with the exception of specificity when using synthetic data with synthetic labels which remained the same. Training on synthetic data using synthetic labels as ground truth alone achieved highest AUC, AP, specificity and PPV among classifiers assessed. CCELLA generated a negative classification by default when text was missing from an image, resulting in a lower proportion of positive labels (Pos) for training for $S_{S}$. All 468 cases in the training dataset where CCELLA extracted negative classification from a positive source image occurred when clinical report text was missing. The test set had a positive case percentage of 56%.
5 Discussion
This study sought to introduce a new framework for training higher-quality synthetic prostate MRI using a limited and fixed volume dataset with limited supervised labeling. This study achieved this by proposing CCELLA and a joint loss function, in addition to reusing pretrained components where possible. Our proposed strategy achieved higher performance than baseline LDMs conditioned on nothing, on radiology label only, and on text only using ELLA. The first ablation study sought to understand the mechanism by which CCELLA achieved better performance. This ablation study trained: 1) an LDM receiving text conditioning by cross-attention using ELLA and timestep conditioning using the ground truth radiology label, 2) an LDM fitted with CCELLA but without using the proposed joint loss. CCELLA outperformed both of these variants, confirming the primary study hypothesis that it is not simply the addition of a class label in the timestep embedding (DECOUPLED) nor the extension of the text adapter for timestep conditioning (L1ONLY) that results in CCELLA’s success. In combination, these ablation results indicate that the joint training of the classifier with the text adapter encourages CCELLA to extract more clinically relevant information from text encoded by a non-clinically-trained LLM. These ablation results also suggest that pretrained, non-medical LLMs could be better adapted for medical text processing when trained jointly with clinical labels. Further investigation should be performed to confirm these findings in different medical domains.
CCELLA was also compared against alternate designs in a second ablation study. Interestingly, these results indicate a near-inverse relationship between diffusion output quality and classifier accuracy, with CCELLA achieving the best in output quality but the worst in classifier accuracy and the SHARED variant achieving the worst in image generation but the best in classifier accuracy. While an initial intuition of these results might suggest that there is a trade-off between the classifier and text adapter performances, further investigation should seek to further explore this relationship and understand how it might be better leveraged for both improved LDM and classifier performance.
Figure 2 shows four sample real images alongside the synthetic images generated by CCELLA and ELLAMAISI using the same corresponding text prompts. Qualitatively, the CCELLA images exhibit natural variance in the sizes and shapes of the prostate while consistently depicting the different anatomical regions within the prostate. CCELLA also did not struggle with image brightness and contrast to the same degree as ELLAMAISI for the same text prompts. ELLAMAISI exhibited occasionally difficulty in defining prostate regions, showing a clear artificial smoothness in Case 4 of Figure 2. The qualitative improvements of CCELLA over ELLAMAISI reaffirm the design choice for the CCELLA adapter to use class conditioning as a directly trainable output.
Finally, this study trained downstream classifier models for prostate cancer prediction from MRI using real data and synthetic data generated by a CCELLA-trained LDM. The $R+S_{R}$ and $R+S_{S}$ classifiers performed equally or better than the $R$ classifier on all metrics, with $R+S_{R}$ outperforming $R$ on all confusion matrix metrics. These results validate the hypothesis that synthetic data can be used to successfully augment the training dataset for a classifier model and aligns with recent literature that generate ground truths alongside data samples. $S_{S}$ also performed comparably to $R$ with higher AUC and AP but lower accuracy, suggesting that the CCELLA-generated synthetic data may be as useful as real data when training downstream machine learning tasks after classifier threshold tuning. These results are promising, though it is important to note that the synthetic data was generated from the same distribution as the real data. Future work should seek to investigate these results further by generating synthetic data for training with real data of different distribution.
Interestingly and in contrast to expectation, the $S_{S}$ classifier outperformed not only the $S_{R}$ classifier in all metrics except sensitivity, but also the $R$ classifier in multiple metrics (AUC, AP, specificity, PPV). Seemingly paradoxically, augmenting the real dataset with synthetic data in $R+S_{R}$ resulted in better performance than augmenting in $R+S_{S}$ in all metrics except AP and sensitivity where they were equal. This likely occurred for three reasons. Firstly, only 3701 of the 5071 training samples had clinical report text. Because null text reports generated negative class predictions by default, adding synthetic data with synthetic labeling in $R+S_{S}$ resulted in a lower positive case percentage than using real labels in $R+S_{R}$ (43% versus 48%). It therefore makes retrospective sense that $R+S_{S}$ would exhibit more difficulty than $R+S_{R}$ in adapting to a test set with higher positive case percentage (56%), resulting in lower accuracy (0.71 versus 0.74). Secondly and related to the first, the training cases without radiology text belonged to both positive and negative classes. Even though the joint classifier loss was removed for these cases, the LDM would still have learned to generate both positive and negative images when presented with a null prompt. This in-turn could have affected the performance of classifiers trained with synthetic data. Thirdly, the CCELLA LDM did not receive the real image as input during inference. However, the ground-truth radiology classification was based on the radiologist re-interpretation of the full image under the most recent prostate MRI reporting standard instead of simply extracting the classification in the clinical text. Consequently, the ground truth re-classification may have lost its meaning in favor of what the LLM interpreted from the clinical text. This could have resulted in the synthetic images aligning well with the text interpretation and CCELLA class extraction but not necessarily with the radiologist re-classification of the real image corresponding to that text, which could explain why $S_{R}$ exhibited poorer performance. Thus, while the same train and test datasets used for LDM training were used for the downstream classifier task for the sake of consistency, future work should seek to further explore the difference between using real versus synthetic labels for synthetic images for the purpose of dataset augmentation when training a model for a medical task on data.
This study has a few limitations, each of which present an opportunity for future work. Firstly, this study emphasized data from a single institution, for a single modality in a single pathology. While the pathology and institution restrictions were intentional to the study design in developing a limited-data LDM training framework, future work should seek to assess method generalizability by adapting this framework: a) to multi-institutional data, b) to images of different modality or pathology. Additionally, prostate MRI is clinically assessed in a minimum of three, and preferably four different MRI sequences [33, 34]. Consequently, portions of the clinical text used for conditioning would have referred to MRI sequences that the LDM was neither trained on nor asked to generate. The use of one sequence alone in this study was chosen to focus primarily on the adapter innovation and data-limited framework design instead of the additional challenge of multi-sequence generation. Nonetheless, future work should seek to extend data-limited LDM training strategies for optimal multi-sequence generation. As well, this study included medical images from public datasets that did not have accompanying text reports for the purpose of gathering more training samples to train the denoising U-Net and the downstream classifier models. While outside the scope of this study, this prompts two opportunities for future work. One opportunity is to investigate how the ratio of conditioned to unconditioned data samples affect model training and whether there is a critical ratio below which training with conditioning becomes unstable or negatively affects training. Another opportunity is to explore synthetic clinical text report generation for these data samples. Finally, the intent of this study aligns with a growing body of literature seeking to use LDMs to generate synthetic medical data for the purpose of mitigating data scarcity challenges within machine learning for medical imaging applications. However, recent literature has also shown that LDMs can occasionally memorize and reproduce their training samples on inference [35]. While outside the scope of this study, future work investigating and improving upon differential privacy within medical imaging LDMs would be valuable.
6 Conclusion
In summary, this study presents CCELLA: a novel adapter for simultaneous alignment of non-medical LLMs to the medical image space and radiology label extraction from medical text for LDM conditioning. This study also presents a CCELLA-centric framework for medical image LDM training given limited data volume and annotation. Using limited data and minimal human annotation, our proposed framework outperforms a recent medical LDM foundation model and successfully augments the training dataset of a downstream classifier model for improved cancer prediction performance over training on real data alone. The results of this study suggest that CCELLA can successfully adapt pretrained, non-medical LLMs for use in conditioning LDMs for medical image generation, and can successfully train a 3D medical image LDM using data generated during clinical routine. Future work should focus on expanding this framework to multi-sequence image generation, validating the framework in a different medical imaging and pathology space, and exploring differential privacy within medical imaging LDMs.